# Combinatorial interpretations of truncated versions of a identity of Gauss

### v2

@@
Keywords: truncated versions, a identity of Gauss, the minimal excludant integer, overpartitions, non-overlined parts
@@ 1 Introduction
A partition $\pi$ of a positive integer $n$ is a finite non-increasing sequence of positive integers $\pi=(\pi_{1},\pi_{2},\ldots,\pi_{m})$ such that $\pi_{1}+\pi_{2}+\cdots+\pi_{m}=n$. The empty sequence forms the only partition of zero. The $\pi_{i}$ are called the parts of $\pi$. Let $p(n)$ denote the number of partitions of $n$. The generating function of $p(n)$ was given by Euler
@@
Apart from Euler’s pentagonal number theorem, there is one other classical theta identity due to Gauss [1, p. 23]:
@@
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+\sum_{j=-k}^{k-1}(-1)^{j}q^{j^{2}}\right)=1+(-1)^{k-1}\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$ (1.9)
Li [12] and Wang and Xiao [15, 16] presented the combinatorial interpretations of the sum in (1.8).
@@ 2 Main results
In this section, we will list the main results in article. One of the main objectives of this article is to present a new combinatorial proof of (1.6).
@@
[18, Definition 4.1] Let $\overline{mex}_{A,a}(\pi)$ be the smallest integer congruent to $a$ modulo $A$ that is not a non-overlined part in the overpartition $\pi$. For $n\geq 1$, define $op_{A,a}(n)$ (resp. $\overline{op}_{A,a}(n)$) to be the number of overpartitions $\pi$ of $n$ such that $\overline{mex}_{A,a}(\pi)$ is congruent to $a$ (resp. $a+A$) modulo $2A$.
@@
$op_{2,1}(n)=op_{2,1}(n,0)\text{ and }\overline{op}_{2,1}(n)=op_{2,1}(n,1).$ (2.2)
In this article, we will present new combinatorial interpretations of the sum in (1.5) and (1.8) by using the minimal excludant non-overlined part of an overpartition.
@@
Combining Theorems 1.3 and 2.2, we can get the following corollary.
@@
$(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k^{2})=\frac{\overline{M}_{k-1}(n)}{2}-\frac{\overline{M}_{k}(n)}{2}.$ (2.6)
Then, we can get (1.7) by using (1.1) with $q\rightarrow q^{\ell}$, (2.4), (2.5), and Euler’s partition identity [9] (see also [1, (1.2.5)]):
@@
Appealing to (1.9) and (2.6), we can get following corollary.
@@
In this article, we will present an analytic proof of Corollary 2.5.
This article is organized as follows. We first give a combinatorial proof of (1.6) in Section 3. Then, we show Theorem 2.2 in Section 4. Finally, we give an analytic proof of Corollary 2.5 in Section 5.
@@
For $n\geq 1$, let $\mathcal{B}(n-1)$ be the set of overpartitions $\lambda$ of n1n-1 such that if $\overline{1}$ appears in $\lambda$ then the smallest part greater than $1$ in $\lambda$ is greater than or equal to two plus the number of parts $1$ in $\lambda$. For example, we have
@@
For $n\geq 1$, we set $\mathcal{C}(n-1)$ be the set of overpartitions of n1n-1 not in $\mathcal{B}(n-1)$, that is, $\mathcal{C}(n-1)$ is the set of overpartitions $\lambda$ of n1n-1 such that $\overline{1}$ appears in $\lambda$ and the smallest part greater than $1$ in $\lambda$ is less than two plus the number of parts $1$ in $\lambda$. For example, we have
@@
4 Proof of Theorem 2.2
The objective of this section is to give a proof of Theorem 2.2. To do this, we need the following lemma.
@@
In [18], Yang and Zhou obtained the following theorem.
Theorem 4.2.
[18, Theorem 4.2] For $n\geq 1$, we have
$op_{2,1}(n)=\overline{p}(n)/2.$
Combining with (2.1) and (2.2), we can get the following lemma. We also give a proof of the following lemma with the aid of Lemma 4.1.
Lemma 4.3.
For $n\geq 1$,
$\overline{p}(n)=2op_{2,1}(n,1).$
Proof. Divide both sides of (1.2) by $\frac{(q;q)_{\infty}}{(-q;q)_{\infty}}$, we can get
@@
Now, we are in a position to give a proof of Theorem 2.2.
Proof of Theorem 2.2. It follows from Lemma 4.1 that
$\overline{p}(n-k^{2})=op_{2,1}(n,k)+op_{2,1}(n,k+1).$
So, we just need to show (2.3). By Lemmas 4.1 and 4.3, we have
$\quad(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)$
$=(-1)^{k}\left(2op_{2,1}(n,1)+2\sum_{j=1}^{k}(-1)^{j}(op_{2,1}(n,j)+op_{2,1}(n,j+1))\right)$
$=(-1)^{k}(-1)^{k}2op_{2,1}(n,k+1)$
$=2op_{2,1}(n,k+1).$
We arrive at (2.3), and thus the proof is complete.
@@
In this section, we aim to give an analytic proof of Corollary 2.5. Note that the generating function of $\overline{M}_{k}(n)$ is

### v3

@@
Keywords: truncated versions, combinatorial interpretations, a identity of Gauss, the minimal excludant integer, overpartitions
@@ 1 Introduction
A partition $\pi$ of a positive integer $n$ is a finite non-increasing sequence of positive integers $\pi=(\pi_{1},\pi_{2},\ldots,\pi_{m})$ such that $\pi_{1}+\pi_{2}+\cdots+\pi_{m}=n$. The empty sequence forms the only partition of zero. The $\pi_{i}$ are called the parts of $\pi$. Let $p(n)$ denote the number of partitions of $n$. The generating function of $p(n)$ was given by Euler.
@@
Apart from Euler’s pentagonal number theorem, there is one other classical theta identity due to Gauss [1, (2.2.12)]:
@@
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\sum_{j=-k}^{k-1}(-1)^{j}q^{j^{2}}=1+(-1)^{k-1}\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}},$ (1.9)
based on which Li [11] presented a combinatorial interpretation of the sum in (1.8).
Theorem 1.6.
[11, Theorem 1.2] For $n,k\geq 1$,
$(-1)^{k-1}\sum_{j=-k}^{k-1}(-1)^{j}\overline{p}(n-j^{2})=\overline{N}_{k}(n),$
where $\overline{N}_{k}(n)$ is the number of overpartitions of $n$ in which a part of size $k$ has to be overlined, the smallest part >k1>k-1 appears exactly $k$ times and it cannot be overlined.
In [14, 15], Wang and Xiao also obtained (1.9) and gave a combinatorial interpretation of the sum in (1.8). Li [11] generalized the sums in (1.5) and (1.8).
Theorem 1.7.
[11, Corollary 5.3] For integers $m$ and $k$ with $m\leq k$,
$(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})\geq 0.$ (1.10)
@@ 2 Main results
In this section, we will list the results obtained in this article. One of the main objectives of this article is to present a new combinatorial proof of (1.6).
@@
[17, Definition 4.1] Let $\overline{mex}_{A,a}(\pi)$ be the smallest positive integer congruent to $a$ modulo $A$ that is not a non-overlined part in the overpartition $\pi$. For $n\geq 1$, define $op_{A,a}(n)$ (resp. $\overline{op}_{A,a}(n)$) to be the number of overpartitions $\pi$ of $n$ such that $\overline{mex}_{A,a}(\pi)$ is congruent to $a$ (resp. $a+A$) modulo $2A$.
@@
Yang and Zhou [17] obtained the following theorem.
Theorem 2.2.
[17, Theorem 4.2] For $n\geq 1$, we have
$op_{2,1}(n)=\overline{p}(n)/{2}.$ (2.2)
@@
$op_{2,1}(n)=op_{2,1}(n,0)\text{ and }\overline{op}_{2,1}(n)=op_{2,1}(n,1).$
Combining with (2.1) and (2.2), we can get the following theorem.
@@
For $n\geq 1$,
$op_{2,1}(n,1)=\overline{p}(n)/{2}.$
In this article, we will give another proof of Theorem 2.3. Then, we will present a combinatorial interpretation of the sum in (1.10) from the point of view of the minimal excludant non-overlined part of an overpartition.
Theorem 2.4.
For integers $m$ and $k$ with $m\leq k$,
$(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=op_{2,1}(n,a)+(-1)^{m+k}op_{2,1}(n,b+1)\text{ if }mk>0,$ (2.3)
and
$(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=op_{2,1}(n,a+1)+(-1)^{m+k}op_{2,1}(n,b+1)\text{ if }mk\leq 0,$ (2.4)
where $a=\text{min}\{|m|,|k|\}$ and $b=\text{max}\{|m|,|k|\}$.
Setting m=k<0m=-k<0 and $m=-k+1\leq 0$ in (2.4), we can get new combinatorial interpretations of the sums in (1.5) and (1.8) respectively.
Corollary 2.5.
@@
Combining Theorem 1.3, Theorem 1.6 and Corollary 2.5, we can get the following corollary.
@@
$op_{2,1}(n,k)-op_{2,1}(n,k+1)=\overline{N}_{k}(n).$
Then, we can get (1.7) by using (1.1) with $q\rightarrow q^{\ell}$, (2.6), (2.7), and Euler’s partition identity [1, (1.2.5)]:
@@
Note that $\overline{M}_{0}(n)=\overline{p}(n)$ for $n\geq 1$, then by Theorem 2.3, we have
$op_{2,1}(n,1)=\overline{M}_{0}(n)/{2}.$
Combining with Corollary 2.7, we get the following result.
@@
This article is organized as follows. We first give a combinatorial proof of (1.6) in Section 3. Then, we show Theorems 2.3 and 2.4 in Section 4. Finally, we give an analytic proof of Corollary 2.8 in Section 5.
@@
For $n\geq 1$, let $\mathcal{B}(n-1)$ be the set of overpartitions $\lambda$ of n1n-1 such that if $\overline{1}$ appears in $\lambda$ then the smallest part lager than $1$ in $\lambda$ is greater than or equal to two plus the number of parts $1$ in $\lambda$. For example, we have
@@
For $n\geq 1$, we set $\mathcal{C}(n-1)$ be the set of overpartitions of n1n-1 not in $\mathcal{B}(n-1)$, that is, $\mathcal{C}(n-1)$ is the set of overpartitions $\lambda$ of n1n-1 such that $\overline{1}$ appears in $\lambda$ and the smallest part larger than $1$ in $\lambda$ is less than two plus the number of parts $1$ in $\lambda$. For example, we have
@@
4 Proofs of Theorems 2.3 and 2.4
The objective of this section is to give the proofs of Theorems 2.3 and 2.4. To do this, we need the following lemma.
@@
Then, we give a proof of Theorem 2.3 with the aid of Lemma 4.1.
Proof of Theorem 2.3. Divide both sides of (1.2) by $\frac{(q;q)_{\infty}}{(-q;q)_{\infty}}$, we can get
@@
We conclude this section with a proof of Theorem 2.4.
Proof of Theorem 2.4. We first show (2.3). There are the following two cases.
Case 1: $k\geq m\geq 1$. In this case, we have $\text{min}\{|m|,k\}=|m|=m$, $\text{min}\{|m|,|k|\}=|m|=m$ and $\text{max}\{|m|,|k|\}=|k|=k$. It follows from Lemma 4.1 that
$\quad(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})$
$=(-1)^{m}\sum_{j=m}^{k}(-1)^{j}\left(op_{2,1}(n,j)+op_{2,1}(n,j+1)\right)$
$=(-1)^{m}\left((-1)^{m}op_{2,1}(n,m)+(-1)^{k}op_{2,1}(n,k+1)\right)$
$=op_{2,1}(n,m)+(-1)^{m+k}op_{2,1}(n,k+1),$
which agrees with (2.3) for $k\geq m\geq 1$.
Case 2: $m\leq k\leq-1$. In this case, we have $\text{min}\{|m|,k\}=k$, $\text{min}\{|m|,|k|\}=|k|=-k$ and $\text{max}\{|m|,|k|\}=|m|=-m$. Again by Lemma 4.1, we have
$\quad(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})$
$=(-1)^{k}\sum_{j=-k}^{-m}(-1)^{j}\overline{p}(n-j^{2})$
$=(-1)^{k}\sum_{j=-k}^{-m}(-1)^{j}\left(op_{2,1}(n,j)+op_{2,1}(n,j+1)\right)$
$=(-1)^{k}\left((-1)^{-k}op_{2,1}(n,-k)+(-1)^{-m}op_{2,1}(n,-m+1)\right)$
$=op_{2,1}(n,-k)+(-1)^{m+k}op_{2,1}(n,-m+1).$
We arrive at (2.3) for $m\leq k\leq-1$.
Now, we have completed the proof of (2.3). Then, we proceed to show (2.4). Under the condition that $m\leq k$ and $mk\leq 0$, we have $m\leq 0\leq k$. Clearly, we have
$\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=\sum_{j=m}^{-1}(-1)^{j}\overline{p}(n-j^{2})+\overline{p}(n)+\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2}).$ (4.2)
In view of (2.3), we get
$\sum_{j=m}^{-1}(-1)^{j}\overline{p}(n-j^{2})=-\left(op_{2,1}(n,1)+(-1)^{m+1}op_{2,1}(n,-m+1)\right),$ (4.3)
and
$\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})=-\left(op_{2,1}(n,1)+(-1)^{k+1}op_{2,1}(n,k+1)\right).$ (4.4)
Substituting (4.3) and (4.4) into (4.2) and combining with Theorem 2.3, we have
$\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=(-1)^{m}op_{2,1}(n,-m+1)+(-1)^{k}op_{2,1}(n,k+1).$ (4.5)
Then, we consider the following two cases.
Case 1: $k\geq-m\geq 0$. In this case, we have $\text{min}\{|m|,k\}=|m|=-m$, $\text{min}\{|m|,|k|\}=|m|=-m$ and $\text{max}\{|m|,|k|\}=|k|=k$. Appealing to (4.5), we get
$\quad(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=op_{2,1}(n,-m+1)+(-1)^{m+k}op_{2,1}(n,k+1).$
So, (2.4) is valid for $k\geq-m\geq 0$.
Case 2: $m>k\geq 0$. In this case, we have $\text{min}\{|m|,k\}=k$, $\text{min}\{|m|,|k|\}=|k|=k$ and $\text{max}\{|m|,|k|\}=|m|=-m$. Using (4.5), we get
$\quad(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=op_{2,1}(n,k+1)+(-1)^{m+k}op_{2,1}(n,-m+1),$
and thus (2.4) is satisfied for $m>k\geq 0$. This completes the proof.
@@
In this section, we aim to give an analytic proof of Corollary 2.8. Note that for $k\geq 0$, the generating function of $\overline{M}_{k}(n)$ is